### v2

# Leveraging Self-Supervised Audio-Visual Pretrained Models to Improve Vocoded Speech Intelligibility in Cochlear Implant Simulation

Abstract
[…]
I Introduction
[…]
Previous studies have confirmed that under quiet conditions, CI can effectively enhance the hearing capability of recipients, especially for speech recognition [97, 96, 29, 28]. However, it is also been reported that speech recognition performance degraded considerably when the target speech signals are distorted [6, 30, 76, 71]. In real-world scenarios, there are several distortion sources, including background noise, reverberation and interfering speech. To address speech distortion issues, a speech enhancement (SE) unit is usually adopted as a front-end processing unit in CI devices [31, 87]. Various techniques, including single-channel SE algorithms like spectral subtraction [85, 46], subspace methods [54], optimized gain functions [59], and commercial solutions [12, 13], have all been applied to improve CI performance. Furthermore, multi-microphone and beamforming approaches have been explored for SE in CI users, taking advantage of spatial filtering to better isolate speech signals from background noise [32, 44, 45, 81].
In recent years, SE techniques have notably improved thanks to advances in machine learning algorithms. Notable examples include non-negative matrix factorization [91, 65], sparse coding [74, 79], compressive sensing [88], and robust principle component analysis [36]. More recently, further enhanced performance has been achieved using the strong regression capabilities of deep learning-based neural networks [55, 92, 86, 93, 98, 17, 34, 89, 94, 68, 69]. For these approaches, deep neural networks are often used as a mapping function to carry out enhancement filtering on noisy input to attain high-quality speech signals. Several extensions have been made to these deep-learning models. One direction is to use a more suitable objective function to train the SE system. In [47, 23, 48, 27, 22, 24, 34], speech metric-oriented objective functions are derived, which can be divided into two categories. The first category directly considers a particular metric to form the objective function, such as [47, 23, 48]. The second uses another neural network model to form the objective function, such as [27, 22, 24, 34]. Experimental results confirm that when a speech-metric oriented objective function is used, the SE system can be guided to achieve desirable output with optimal speech metric scores. In addition to designing more suitable objective functions, some researchers have attempted to incorporate information from other modalities as auxiliary inputs to the SE model, enabling exploitation of additional contextual information. Visual clues are one important modality that carries complementary information to speech signals during everyday communication. For example, the McGurk effect [61] demonstrates the cross-effect between the visualized mouth or lip shape and human hearing perception. From this viewpoint, numerous audio-visual multi-modal SE approaches, termed AVSE, have been proposed [25, 16, 33, 62, 41, 73, 63, 9, 8]. These studies clearly show that visual cues can successfully enhance the performance of audio-only speech enhancement (AOSE).
Developing an efficient AVSE system with limited training data is a crucial undertaking in real-world scenarios. Following [8], we address this challenge by exploiting self-supervised learning (SSL) in AVSE. The motivation for our work is the ability of SSL models to automatically label features and generate them based on knowledge of the input data. Specifically, instead of providing simple labels for each object as is the case for most supervised methods, SSL allows the extraction of more representative features such as the relative position of certain objects and their contextual relationship. This boosts performance by improving inference and reduces the amount of raw data required for pre-training [49, 14]. For example, the well-known Bidirectional Encoder Representations from Transformers (BERT) model produces contextual representations of languages generated using only a plain text corpus. BERT is a versatile AI pre-training model originally developed for Natural Language Processing (NLP). Compared to previous models using supervised methods, BERT has been shown to deliver substantial improvements in performance for a wide variety of tasks such as NLP and word recognition [43, 78, 37]. HuBERT, a BERT-derived model using hidden-units [35], has proven to be effective for denoising from an noise speech [38, 39]. In this paper, we develop a novel SSL-AVSE framework by leveraging AV-HuBERT [77], which is a HuBERT-based audio-visual model.
In the past decade, deep learning-based SE techniques have been applied to CI systems [4, 50, 51, 42, 58], enabling more adaptive and robust processing methods tailored to the diverse auditory environments experienced by CI users. While the benefits of incorporating visual cues into the SE process are well-established, their specific advantages for CI devices have, to our knowledge, not yet been explored. In this study, we aim to evaluate the effectiveness of the proposed SSL-AVSE system for CI, with the implementation and user test scenario illustrated in Fig. 1. As shown in the figure, the SSL-AVSE system processes the combined noisy speech and video as inputs and outputs enhanced speech, which is then provided to a CI device. In this study, we employ a 16-channel speech vocoder to process enhanced utterances, simulating CI audio and calculating relevant performance metrics. Additionally, a listening test with 80 subjects is conducted to assess subjective speech quality and word intelligibility. Results indicate that our SSL-AVSE model achieves a maximum improvement of over 50% and 45% in speech quality and intelligibility compared to the original noisy signals, respectively, and a maximum improvement of over 20% and 40% compared to the AOSE system, respectively.
Figure 1: Overview of our proposed system: Noisy speech and video are input into our SSL-based AVSE model, which outputs enhanced speech for CI users.
II Methods
In this section, we first formulate our problem and then present the SSL-AVSE network used in this study. We then demonstrate our training criteria and inference procedures, including the mathematical theories involved.
II-A Problem Formulation
Suppose we have a speech signal $s(t)$ and a noise signal $n(t)$. The noisy signal $x(t)$ can then be denoted as:
$x(t)=s(t)+n(t)$ (1)
Next, the noisy signal $x(t)$ is transformed into the spectral feature $X$. The model predicts a ratio mask $M$ to extract the clean speech signal for the corresponding target speaker from $X$. $M$ is trained from our SSL-AVSE model, which uses the noisy signal $x(t)$ and additional visual cues $v(t)$ which is comprised of the lip image sequence. The enhanced speech can then be obtained by the following formula:
$\hat{S}=X\otimes M$ (2)
where $\otimes$” indicates an element-wise multiplication. From the above, we can then use the enhanced representation, $\hat{S}$, to reconstruct the waveform signal $\hat{s}(t)$, also known as enhanced audio.
II-B Audio-Visual Speech Enhancement Networks
In this paper, we propose a novel SSL-AVSE framework, as shown in Fig. 2. Specifically, we combine the SE model with a Transformer-based AV-HuBERT network using self-attention mechanisms to provide enhanced performance.
II-B1 Data Preprocessing
The model takes the visual stream of detected lip images from the target speaker $v(t)$, the noisy speech $x(t)$ as input and outputs the enhanced speech $\hat{s}(t)$ for the target speaker while suppressing noise sources.
Audio Preprocessing. In this study, the noisy speech signal $x(t)$ is first transformed into spectrograms using the Short-Time Fourier Transform (STFT). Subsequently, the $log1p$ function ($log1p(z)$ = log(1+$z$)) is applied to these spectrograms to extract $log1p$ spectral features. It has been demonstrated in our prior research that these $log1p$ spectral features outperform conventional log power spectral features in terms of SE performance [39, 21]. As shown in Fig. 2, the noisy $log1p$ spectral features are multiplied by the estimated mask to produce enhanced speech, as shown in Eq. (2). The mask is generated by our SSL-AVSE system, which is based on AV-Hubert and takes the video signal, $v(t)$, along with the noisy speech signal, $x(t)$, as input.
Video Preprocessing. The video signal, $v(t)$, is comprised of the sequential images sampled at 25 fps and are cropped around the target speaker’s mouth region of interest (ROI). This is determined with a facial landmark detector consisting of a two-dimensional Face Alignment Network (FAN)[5]. Assuming a 3-second video clip at 25 fps, we can obtain a total of 75 lip images from the target speaker.
II-B2 Model
The representations of each transformer-encoder layer are denoted as $H^{l}$, where $10\leq l\leq L-$ 1 and $L$ is the number of layers. A trainable function $w(\cdot)$ is then applied to all of the layer representations as follows:
$H_{WS}=\sum_{l=0}^{L-1}w^{l}H^{l},$ (3)
where $w^{l}$ is the weight of the $l$-th layer and has the properties $w^{l}\geq 0$ and $\sum_{l}w^{l}=1$.
The features are then fed into the SE model consisting of a two-layer bidirectional long short-term memory (BLSTM) module sandwiched between two linear layers which performs forward and back propagation simultaneously. The output of the SE module is a soft mask and is multiplied by the magnitude of the noisy speech spectra. This is then compared with the clean speech spectra to determine the L1 (absolute) loss.
While the video segment length used for learning SSL-AVSE is fixed, at the inference stage, our audiovisual extraction model can be applied to process videos of arbitrary length. This is done by applying a sliding window technique, which shifts the proposed window along the video segment until its entire length is covered.
Figure 2: The proposed SSL-AVSE model includes AV-HuBERT and SE modules. The AV-HuBERT model consisting of multiple transformer layers is used to extract key features from noisy speech and lip images. An SE model then performs enhancement using them.
II-C CI Vocoded Speech
We passed voice signals through vocoder to simulate CI sounds. These simulations were then played to normal hearing (NH) people to conduct an audiometric test [18, 75]. Compared with ordinary speech, vocoded speech is more difficult to understand by NH listeners due to the loss of spectral detail. Several studies have examined CI vocoder simulated speech on NH subjects in order to understand the associations between specific factors and CI users[80, 20, 19, 50]. Because accurate CI sounds are not always readily available, vocoder simulations can avoid the manifestation of patient-specific confounding factors, such as neural survival patterns [53]. Therefore, the CI vocoder can serve as an invaluable tool in related research.
To simulate CI audio in this study, the tone vocoder was performed for an audio signal by using the procedure depicted in Fig. 3. As shown in the figure, there are four steps: (1) 16 Butterworth band-pass filters were used to process an input temporal sequence to produce bandpass signals. (2) For each band waveform, a full-wave rectification function was leveraged to smooth the signal and to generate the corresponding envelope wave. (3) We added a tonal signal to the envelope to produce the modulated band voice. (4) We summed all modulated voice and performed normalization operation to generate the vocoded speech, which has the identical root-mean-square value to the original input signal.
Figure 3: Block diagram of our tone vocoder. The system consists of three modules: (1) A band-pass filter, (2) An envelope extractor and (3) a compressor. The results of each band are then summed and combined to form vocoded speech.
II-D Measures of Speech Intelligibility
Speech intelligibility is assessed through objective and subjective measures in scientific research. The three most widely used measures include the Perceptual Evaluation of Speech Quality (PESQ) [70] and the Short Time Objective Intelligibility (STOI) [82] for non-vocoded speech and the Normalized Covariance Metric (NCM) [57] for vocoded speech. PESQ is an algorithm designed to predict subjective opinion scores of a degraded audio sample. It provides numerical values ranging from -0.5 to 4.5. This measure is considered a full reference algorithm as it compares both the original and degraded signal [84]. STOI is another important measure of speech quality that is often used in SE settings. It is highly correlated with the intelligibility of degraded speech signals and takes into consideration the effect of non-linear processing on noisy speech. NCM is derived from multivariate normal distribution and is used to determine the correlation between multiple variables which are usually described as a matrix. Its dimension is dependent on the number of channels involved; since our vocoder simulator consists of 16 distinctive channels, a 16-dimensional NCM can be used to calculate their correlation coefficients. This value depends on the relationship between enhanced and clean speech; the greater the similarity the larger the NCM value. Orthogonal matrices, which consist of completely independent variables, have a correlation coefficient of 0, while congruent matrices, which consist of completely dependent variables, have a correlation coefficient of 1. To compare performance, we used PESQ and STOI scores to evaluate the SE results for normal speech. For CI-vocoded speech, we used NCM for performance comparison, as suggested in [7].
Subjective measures used in this study include overall speech quality and word intelligibility. The former is determined when subjects are asked to rate the quality of an entire utterance from 1 to 5. The higher the score, the better the perceived quality of the recorded sentence. The latter is based on whether subjects can clearly understand individual words in the given sentence. Since each utterance contains a total of 10 Chinese characters, a score of 1.0 (=10/10) indicates a complete understanding of each character in the given target sentence.
II-E Implementation Details
II-E1 Ablation Studies
To confirm the effectiveness of incorporating the pre-trained AV model, we built an AOSE system employing a BLSTM (Bidirectional Long-short Term Memory) model. In addition, we added ResNET-18 to process visual cues and combined it with the original one to form an AVSE model. Finally, we replace ResNET-18 with AV-HuBERT which allows for self-supervised learning.
II-E2 Model Training
We optimized the SSL-AVSE model using partial fine-tuning (PF). Here, the weights pertaining to feature extraction in AV-HuBERT are fixed, whereas the weights in the SSL block are updated based on the pre-trained checkpoint. Training was conducted using an Adam optimizer with a weight decay of $10^{-4}$, a batchsize of 32 and an initial learning rate of $10^{-4}$. In addition, the learning rate is halved when errors are encountered. The proposed model was trained on the TMSV dataset for 50 epochs. We avoid overfitting on the training set by employing early-stopping techniques.
III Experiments
III-A Experimental Setup
Noise can be categorized into static and dynamic based on the acoustic properties in the frequency domain. For example, monotonic background noises would fall under the category of static noise, while highly varied ones would be considered as dynamic. The removing of dynamic noise poses a greater challenge due to the similarities between the input and noise signals such as frequency bands and tonal pitch. To demonstrate the effectiveness of our models, we selected five different dynamic noise types: “babycry” (sound of a baby crying), “babble” (multiple people talking simultaneously in a crowd), “one talker”, “two talkers” and “three talkers”. All these types are associated with human sounds or utterances. The amount of noise used to corrupt the input signal is measured by the signal-to-noise ratio (SNR) and expressed in decibels (dB). The model was trained on the TMSV dataset with video recordings of 18 native Mandarin speakers (13 male and 5 female), each containing 320 utterances. We used 8 speakers (4 male and 4 female) for training and 2 speakers for testing (1 male and 1 female). Each sentence consists of 10 Chinese characters and has a duration of 2 to 4 seconds. The video was recorded with a resolution of 1,920 pixels $\times$ 1,080 pixels at 50 frames per second (fps) while the audio was recorded at a sampling rate of 48 kHz. The first 200 sentences were used to form the training set, while the last 120 sentences were used to form the test sets. To form clean-noisy speech pairs, the training utterances were artificially corrupted by 100 types of noise at five different SNR levels ranging from -12 to 12 dB with an increment of 6 dB. This process generates approximately 600-hour noisy utterances, including the 9-hour randomly selected training data. Instead of using all generated utterances, the training set used in this study comprised 12,000 utterances. Testing was conducted on clean speech mixed with the five dynamic noise types at SNRs ranging from -7 dB to 8 dB at increments of 3 dB.
III-B Experimental Results
III-B1 Effect of Fine-tuning the SSL Pretrained Model
First, we present the speech quality (in terms of PESQ) and speech intelligibility (in terms of STOI) of SSL-AVSE with different pre-trained AV-HuBERT models and verify the effectiveness of fine-tuning these models for AVSE. Table I shows the PESQ and STOI scores of SSL-AVSE using several AV-HuBERT models pre-trained on different datasets, including LRS3 [1] and VoxCeleb2 [10], provided by the authors in [35]. From the table, we can observe that fine-tuning a pre-trained AV-HuBERT model with more diverse data leads to enhanced results. This confirms the effectiveness of our method for leveraging languages with high resources to a use case targeted for another language with low or limited resources. Besides, the results in Table I confirm the efficacy of fine-tuning an AV-HuBERT in our method. In the following experiments, SSL-AVSE denotes the one fine-tuning the AV-HuBERT pre-trained on LRS3, VoxCeleb2 and noise augmentation, which is the best setup.
TABLE I: Objective scores of the speech enhanced by fine-tuning pre-trained AV-HuBERT models. L, V, and N represent LRS3, VoxCeleb2, and noise augmentation, respectively.
PESQ STOI
Noisy 1.434 0.695
SSL-AVSE (L/V/N. w/o fine-tuning) 1.565 0.719
SSL-AVSE (L) 1.615 0.728
SSL-AVSE (L/V) 1.651 0.737
SSL-AVSE (L/V/N) 1.665 0.738
III-B2 Objective Results
TABLE II: Objective PESQ scores for non-vocoded speech. We can see that the lower the SNR, the greater the difference between the PESQ scores of SSL-AVSE enhanced audio and those enhanced by either AVSE or AOSE. The converse is true for higher SNRs; the amount of improvement is less marked. However, the higher SNRs, the greater the difference between the PESQ scores of SSL-AVSE enhanced audio and those of the noisy baseline. A stands for audio-only, while AV stands for audio-visual.
modality -7dB -4dB -1dB 2dB 5dB 8dB
Noisy N/A 1.214 1.260 1.344 1.449 1.586 1.771
AOSE A 1.227 1.296 1.403 1.540 1.713 1.934
ConvTasNet [56] A 1.242 1.311 1.424 1.569 1.764 2.001
VisualVoice [26] AV 1.283 1.349 1.416 1.486 1.667 1.792
AVSE-VAE [72] AV 1.217 1.285 1.388 1.511 1.677 1.890
AVSE AV 1.226 1.324 1.453 1.591 1.773 2.002
SSL-AVSE AV 1.271 1.353 1.474 1.619 1.801 2.020
TABLE III: Objective STOI scores for non-vocoded speech. We can see that the lower the SNR, the greater the difference between the STOI scores of SSL-AVSE enhanced audio and those enhanced by either AVSE or AOSE. For higher SNRs, the results for different enhancement methods converge. A stands for audio-only, while AV stands for audio-visual.
modality -7dB -4dB -1dB 2dB 5dB 8dB
Noisy N/A 0.536 0.598 0.644 0.731 0.794 0.848
AOSE A 0.523 0.590 0.663 0.733 0.797 0.854
ConvTasNet [56] A 0.549 0.611 0.682 0.753 0.808 0.859
VisualVoice [26] AV 0.591 0.635 0.687 0.729 0.767 0.810
AVSE-VAE [72] AV 0.516 0.580 0.647 0.714 0.774 0.827
AVSE AV 0.549 0.616 0.685 0.751 0.809 0.860
SSL-AVSE AV 0.582 0.636 0.695 0.753 0.809 0.858
TABLE IV: Objective NCM scores for vocoded speech. Similar to the results of STOI scores, We can see that the lower the SNR, the greater the difference between the NCM scores of SSL-AVSE enhanced audio and those enhanced by either AVSE or AOSE. However, unlike the STOI results, there is no marked convergence between the absolute amount of improvement between the NCM scores of the noisy baseline and that of utterances enhanced by SSL-AVSE.
-7dB -4dB -1dB 2dB 5dB 8dB
Noisy 0.226 0.321 0.416 0.519 0.600 0.671
AOSE 0.354 0.441 0.553 0.630 0.724 0.811
AVSE 0.387 0.482 0.581 0.681 0.773 0.849
SSL-AVSE 0.423 0.507 0.598 0.690 0.777 0.846
As shown in Tables II to IV, compared with AOSE and AVSE, enhancement by SSL-AVSE resulted in higher objective measures of speech intelligibility for most types of noise. The difference was particularly notable for low SNRs. For a SNR of -7 dB, the PESQ, STOI and NCM values of SSL-AVSE were 3.7% (from 1.226 to 1.271), 6.0% (from 0.547 to 0.582) and 9.3% (from 0.387 to 0.423) higher than those of AVSE, respectively, while they were 3.6% (from 1.227 to 1.271), 11.3% (from 0.523 to 0.582) and 19.5% (from 0.354 to 0.423) higher than those of AOSE, respectively. The results were even more striking when compared with the noisy baseline; PESQ, STOI and NCM values increased by 4.7% (from 1.214 to 1.271), 8.6% (from 0.536 to 0.582) and 87.2% (from 0.226 to 0.423), respectively. However, for higher SNRs, the differences between the PESQ, STOI and NCM values of SSL-AVSE and AVSE enhanced speech are less pronounced; this is probably because automatic labelling is less useful when fewer distinguishing features are present. The presence of visual cues were also found to be less effective at audio denoising for higher SNRs; results also show reduced improvement in all three speech quality scores in audio enhanced by SSL-AVSE over those enhanced by AOSE.
TABLE V: Paired t-test for speech quality under selected noise conditions. The degree of freedom for all values is 19. A p-value of less than 0.05 implies statistical significance.
babble 2 dB babble 5 dB babycry 2 dB babycry 5 dB
t-value $p$-value t-value $p$-value t-value $p$-value t-value $p$-value
AOSE 1.571 $<$0.05 1.746 $<$0.05 1.191 $<$0.05 3.145 $<$0.05
AVSE 1.337 $<$0.01 0.883 $<$0.01 1.047 $<$0.01 2.732 $<$0.05
SSL-AVSE 2.743 $<$0.001 1.402 $<$0.001 1.031 $<$0.001 1.433 $<$0.01
TABLE VI: Paired t-test for speech intelligibility under selected noise conditions. The degree of freedom for all values is 19. A p-value of less than 0.05 implies statistical significance.
babble 2 dB babble 5 dB babycry 2 dB babycry 5 dB
t-value $p$-value t-value $p$-value t-value $p$-value t-value $p$-value
AOSE 2.518 $<$0.05 3.452 $<$0.05 2.346 $<$0.05 2.128 $<$0.05
AVSE 1.664 $<$0.01 3.683 $<$0.05 2.038 $<$0.05 3.038 $<$0.01
SSL-AVSE 2.842 $<$0.001 4.535 $<$0.01 1.406 $<$0.01 4.406 $<$0.01
III-B3 Spectrogram Analysis
A spectrogram plot is frequently employed to visually represents the time-frequency characteristics of a speech signal. In Fig. 4, we present spectra of a noisy speech signal at a 2 dB SNR, enhanced using four methods: AOSE, ConvTasNet, AVSE, and SSL-AVSE. Additionally, the spectra of the corresponding clean speech is included for comparison. It can be observed that SSL-AVSE outperforms AOSE, ConvTasNet and AVSE in effectively removing noise components and restoring the original clean speech. In Fig. 5, we showcase spectra of vocoded speech. From the figure, it is evident that the spectra of vocoded speech processed by SSL-AVSE preserves much clearer speech structures compared with AOSE and AVSE.
Figure 4: Spectrograms of clean, noisy, AOSE enhanced, AVSE enhanced, ConvTasNeT enhanced and SSL-AVSE enhanced audio signals for “babble 2 dB” noise. Note that SSL-AVSE enhanced audio emphasizes audio tones within the range of human speech more than those of other enhancement methods.
Figure 5: Spectrograms of vocoded speech for noisy, clean, AOSE enhanced, AVSE enhanced and SSL-AVSE enhanced audio signals for “babble 2 dB” noise. Note that SSL-AVSE-enhanced speech has clearer structures compared with those of other enhancement methods.
III-B4 Subjective Results
We conducted listening tests using four different types of noises. Each noise types consisted 120 utterances, with 20 from each of the following six categories: AOSE, AVSE, SSL-AVSE, clean, logarithm minimum mean squared error (logMMSE), and noisy speech. The logMMSE method performs SE based on assumptions of distribution and properties of speech and noise signals[15]. Similar to[50], we also include logMMSE as part of our benchmark. In this study, we also selected to present those utterances corrupted by “babycry” and “babble” noises with SNRs of 2 dB and 5 dB, respectively. Since that our objective is to maximize SE in CI devices, we conducted our experiment using vocoded speech. A total of 80 participants with a mean age of 35.7 took part in the study, with 20 individuals assigned to each noise or sound condition to reduce cross-referencing bias. The listening tests were conducted in a quiet room, while the utterances were uploaded onto a specially designed system that presented them randomly to the users.
As shown in Fig. 6 and Fig. 7, results show that for both speech quality and word intelligibility, SSL-AVSE outperformed both AVSE and AOSE. For speech quality scores, when subjected to the most challenging “babble 2 dB” noise condition, the SSL-AVSE model exhibited a 26.0% increase (from 2.42 to 3.05) for the former and a 41.9% increase (from 2.15 to 3.05 ) for the latter, respectively. For word intelligibility scores, the improvement over the former and the latter were 45.6% (from 0.388 to 0.565) and 65.2% (from 0.342 to 0.565), respectively.
Figure 6: Subjective sound quality scores for vocoded speech enhanced with different models. The x-axis represents the noise type while the y-axis represents the sound quality score. For all noise types, SSL-AVSE is shown to perform notably better than other methods, with the greatest improvements occurring for noises with an SNR of 2 dB. logMMSE, a statistical SE method, is shown to perform worse than the noisy baseline.
Figure 7: Subjective character intelligibility scores of speech enhanced using different models. The x-axis represents the noise type while the y-axis represents the word intelligibility score. As in the case of speech quality, SSL-AVSE is also shown to perform notably better than other methods, with the greatest improvements occurring for noises with an SNR of 2 dB.
Figure 8: Relationship between the subjective and objective scores of speech enhanced using different models. We choose utterances corrupted with “babble 2 dB” noise. The x-axis represents the subjective score while the y-axis represents the objective score. The centers of the ovals represent the mean objective and subjective scores, while the lengths of the major and minor axes represent the standard deviations of the objective and subjective scores, respectively. There is clear segregation between utterances enhanced using the three different methods, with SSL-AVSE having the best effect.
We further conducted paired t-tests on the quality and intelligibility results of listening test under SSL-AVSE, AVSE and AOSE conditions to verify their statistical significance. We specifically compare the outcomes obtained from “babble” and “babycry” noises for SNRs of 2dB and 5dB, since they are generally considered the most difficult to denoise. Similar trends are observed for other types of noise as well. Results shown in Tables V to VI indicate that audio enhanced by all three methods differ statistically from noisy audio, with $p$-values all less than 0.001 for “babble” noises and $p$-values all less than 0.05 for “babycry” noises. Since the “babble” noise condition results in smaller $p$-values than those of “babycry”, this demonstrates that our model actually performs better under more challenging noise conditions, especially those with multiple talkers. Further comparisons between audio enhanced by SSL-AVSE and those enhanced by either AVSE or AOSE also show statistical differences, further confirming the utility of our proposed models.
III-B5 Comparison between Objective and Subjective Results
Both objective and subjective evaluations demonstrate that the performance measures improve with SNR levels. Moreover, it is observed that audio signals with lower SNRs exhibit a more substantial enhancement in speech quality, highlighting the robustness of SSL-AVSE in challenging noisy environments. The subjective results further validate the effectiveness of our model in handling vocoded speech, showing improvements comparable to those determined by objective methods when compared to the noisy baseline. To more explicitly explore the relationship between objective and subjective results, three panels are presented in Fig. 8, focusing on utterances corrupted with “babble 2 dB” noise. For all three panels, the x-axis represents the subjective score, while the y-axis represents the objective score. Analyzing the left panel of Fig. 8, it is evident that SSL-AVSE outperforms AVSE and AOSE notably in terms of both objective and subjective speech quality scores, as indicated by the score distributions located towards the top-right side. Furthermore, examining the center and right panels of Fig. 8, it is evident that SSL-AVSE achieves the best performance for both objective and subjective speech intelligibility scores, as indicated by the score distributions also located towards the top-right side.
IV Discussion
In this paper, we propose a deep neural network-based AVSE method that leverages both audio and visual signals, evaluating its effectiveness in enhancing speech perception in CI simulations. As shown in Table IV, when combined with self-supervised learning, NCM values of enhanced audio increased by between 26.5% (from 0.671 to 0.846 for 8dB) and 87.2% (from 0.226 to 0.423 for -7dB) compared with the noisy baseline. This demonstrates the effectiveness of our model on vocoded audio signals. Subjective results also showed improvements of between 12.3% and 45.2% for sound quality (from 3.59 to 4.30 for “babycry 5dB” and from 2.10 to 3.05 for “babble 2dB”) and improvements of between 5.6% and 51.8% for word intelligibility (from 0.654 to 0.691 for “babycry 2dB” and from 0.372 to 0.565 for “babble 2dB”) when compared with the noisy baseline, confirming our results.
Due to the relatively large size of video files, real-time AVSE using both video and audio signals remains a challenging task. First, an appropriate image extraction rate needs to be determined based on trade-offs between testing time and accuracy. Second, most SSL models are relatively large; for example, our base AV-HuBERT model contains more than 300 million parameters. Reducing the number of parameters through prudent data pre-processing or by pruning the model to reduce its size after training yield faster results at the expense of testing accuracy. For settings that require high fidelity output audio including lectures and meetings, this may not be feasible. Further work is required to determine the optimal parameters for each setting. While our current model may not directly suit CI devices with limited capacity, it can be implemented on external interface hardware like smartphones, providing ample computational resources for speech signal processing. For example, CIs have started to connect with phones as a “router” such as in the Cochlear Wireless PhoneClip, which may improve speech recognition by converting Bluetooth sound streams to the Nucleus 6 (CP910) sound processor [40]. Similarly, the AudioLink device can help with auditory training on speech recognition with compatible MED-EL sound processors [66]. More straightforward designs, such as the Nucleus 7 (CP1000) and more recent sound processors may directly connect to compatible Apple (Made for iPhone, MFi) and Android (Android Streaming for Hearing Aids, ASHA) devices [83]. Our system is also well-suited for applications like teleconferencing and playback applications, where real-time processing is not essential. With rapid advancements in hardware, complex systems based on DNNs can now run on edge devices, suggesting that hardware capable of supporting our system will soon be widely available.
V Conclusion
We propose a novel AVSE model that leverages a pretrained audio-visual foundation model to enhance speech intelligibility and quality in CI scenarios. Although the model was pretrained using an English corpus, it performed well on SE tasks involving Mandarin datasets, demonstrating the ability to generalize to different target languages. Compared with other state-of-the-art AVSE methods, our proposed model resulted in a notable increase in speech quality and word intelligibility. In the future, we aim to conduct real-time signal processing by compressing our model using pruning and data quantization techniques. This enables more rapid processing of audio-visual signals without compromising significantly on quality. The results can then be output as audio signals or as text files for the patient, enabling them to regain near-normal hearing function.